OpenStudy (anonymous):

1 find the least number between 200 and 500 which leaves a reminder of 3 in each case when devided by 8,10,12 and 30

OpenStudy (superhelp101):

steps 1. 8 = \[2^{3}\], 10 = 2*5, 12 = 2^2*3, 18 = 2 *3^2 the all divide 2^3 *3^2*5 = 360 363 is least number in 200 to 500 there is a remainder 3 in each case 3 +360 = 363

OpenStudy (superhelp101):

got me? :)

OpenStudy (ikram002p):

let our number be G G mod 8=3 G mod 10=3 G mod 12 =3 G mod 30 = 3